OpenStudy (claire.bracken):

A roll of quarters contains 40 quarters and a roll of dimes contains 50 dimes. James has a jar that contains 83 quarters and 159 dimes. Lindsay has a jar that contains 129 quarters and 266 dimes. James and Lindsay pool these quarters and dimes and make complete rolls with as many of the coins as possible. In dollars, how much are the leftover quarters and dimes worth?

OpenStudy (s4sensitiveandshy):

the question is to find out leftover money Lindsay's jar contains 129 Quarters and other jar contains 83 quarters add them 129+83= 212 each roll of quarters contains 40 quarters so divide 212 by 40 that will give you 5.3 that represents 5 roll of quarters and 3 quarters leftover same way to find dimes can you do that ?

OpenStudy (s4sensitiveandshy):

hmm wait there is a typo

OpenStudy (claire.bracken):

@S4Sensitiveandshy I tried doing that I got 205 what i did was i realized the maximum # of rolls was 5 212-200 = 12 425-250 = 175 12*25 + 1750 = 2050 = $205 but Alcumus on AopS said it was incorrect...

OpenStudy (s4sensitiveandshy):

because each quarter worth .25 not 25 ;)

OpenStudy (claire.bracken):

OH I SEE 12*0.25 + 175 = 178 would that be the answer? @S4Sensitiveandshy

OpenStudy (s4sensitiveandshy):

in my first I meant to say 3 dollar is leftover not quarters 212 subtract 12 200 is divisible by 40 so I get 5 (roll of quarters ) and 12 quarters 12 times .25 = 3 dollar

OpenStudy (s4sensitiveandshy):

do you mean 178 leftover money ?? :o

OpenStudy (claire.bracken):

yup...

OpenStudy (s4sensitiveandshy):

hmmm no that's not correct the total money is not even equal to 178 each dime worth .10 cents only

OpenStudy (claire.bracken):

then 20.5?

OpenStudy (s4sensitiveandshy):

can you please show how did you get 20.5 ?

OpenStudy (claire.bracken):

212-200 = 12 425-250 = 175 12*0.25 + 175*0.1 = 20.5

OpenStudy (s4sensitiveandshy):

why did you subtract 250 from 425 ?? :o

OpenStudy (claire.bracken):

because there needs to be 5 rolls total at max... since 40 can only go into 212, 5 times. (40*5 = 200) and then so that means 50 (the dimes)*5 = 250 dimes... the total number of dimes is 425 so 425-250 dimes = 175 dimes

OpenStudy (s4sensitiveandshy):

hmm but 425-250=175 175 is 100+75 where 100 is also divisible by 50 right ? the simple and easy way is to divide 425 by 50 or subtract 425-25=400 400 is also divisible by 50 so the left over is 25 dimes

OpenStudy (claire.bracken):

OHHH i read it wrong i apologize haha i was thinking of it as 40 quarters and 50 dimes make a roll, not separately

OpenStudy (s4sensitiveandshy):

ohh I see you can't put dime into quarter roll ;) that would be cheating

OpenStudy (claire.bracken):

so the answer is $55 correct?? 12*25 + 25*10 = 550 cents = 55 dollars

OpenStudy (s4sensitiveandshy):

okay if I give you 12 quarters and 25 dimes are you gonna be happy would that worth 55 dollars ?? ;) :P

OpenStudy (s4sensitiveandshy):

wait a sec

OpenStudy (claire.bracken):

oops 5.50

OpenStudy (claire.bracken):

5.50 is right :))) THANKS FOR ALL YOUR HELP!!!

OpenStudy (s4sensitiveandshy):

looks good :D 12 quarters that would be $3 and 25 dimes that would be 2 dollar and 50 cents and then add 3+2.50 would be 5 dollar and .50 cents

OpenStudy (s4sensitiveandshy):

haha you're welcome :)